Decide which of the two points lies on the graph of the line.
2x + y = 10
a (4, 3) b (-4, 18)
1
Expert's answer
2013-01-11T06:35:26-0500
2x + y = 10
a (4, 3) b (-4, 18)
if point (x1, y1) lie on the graph 2x + y = 10 then
2x1+y1=10
1) a (4, 3)
2*4 + 3 = 11 - not equal10
point a does not lies on the line
2) b (-4, 18)
2*4 - 18 = 10 equal 10
point b lies on the line